OpenStudy (anonymous):

I need help with one more problem if someone will help me. I'll put it below:

OpenStudy (anonymous):

\[(2\sqrt{5}+3\sqrt{7})^{2}\]

OpenStudy (anonymous):

(a+b)^2 = (a+b)(a+b)

OpenStudy (anonymous):

So: \[(2\sqrt{5} + 3\sqrt{7})(2\sqrt{5} + 3\sqrt{7})\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

foil

OpenStudy (anonymous):

Thanks @optiquest !

OpenStudy (anonymous):

@optiquest So I got this:

OpenStudy (anonymous):

Ugh it messed up. It was supposed to say\[2\sqrt{5}(2\sqrt{5}+3\sqrt{7})+3\sqrt{7}(2\sqrt{5} + 3\sqrt{7})\]

OpenStudy (anonymous):

I dont think you foiled right

OpenStudy (anonymous):

Yeah I do it differently than my teacher does. I usually get the right answer, but I'm not sure now. :/

OpenStudy (anonymous):

\[(2\sqrt{5})^2 + 3\sqrt{7}\times2\sqrt{5} + (3\sqrt{7})^2\]

OpenStudy (anonymous):

\[2^{2}\sqrt{5}^{2}+3\times7^{1/2}\times2\times5^{1/2}+3^{2}\sqrt{7}^{2}\]

OpenStudy (anonymous):

\[4\times5+3\times 7^{.5}\times2\times5^{.5}+9\times7\]

OpenStudy (anonymous):

see above

OpenStudy (anonymous):

Where did all of those numbers come from?

OpenStudy (anonymous):

\[20+63+6\times(7*5)^{.5}\]

OpenStudy (anonymous):

which ones

OpenStudy (anonymous):

The 5 and 3 and 9

OpenStudy (anonymous):

im not sure where you are looking

OpenStudy (anonymous):

Okay I understand every part except for the exponents. What is the .5 from?

OpenStudy (anonymous):

you can just rewrite a root at a number raised to 1/2 or .5 they are equivalent

OpenStudy (anonymous):

@Hero do think you could help me out?

OpenStudy (anonymous):

\[x ^{.5} = x ^{1/2} = \sqrt{x}\]

OpenStudy (amistre64):

\[2\sqrt{5}(2\sqrt{5}+3\sqrt{7})+3\sqrt{7}(2\sqrt{5} + 3\sqrt{7})\] youre distribution is fine, notice we have 2 more distributions, finish it out \[2\sqrt{5}(2\sqrt{5})+2\sqrt{5}(3\sqrt{7})+3\sqrt{7}(2\sqrt{5}) + 3\sqrt{7}(3\sqrt{7})\] \[4\sqrt{25}+6\sqrt{35}+6\sqrt{35} + 9\sqrt{49}\]